# TIF-UNIMI-2025-23

### v1

@@ 7 The resummed coefficient function
The resummed result Eq. (6.3) provides an expression for the physical anomalous dimension, i.e. essentially for the scale dependence of the coefficient function. Hence, a little more work is required in order to arrive at a resummed expression for the coefficient function itself. This is done by separating off the part of the $Q^{2}$ dependence of the coefficient function that goes through the ratio to the factorization scale $Q^{2}/\mu_{F}^{2}$. The final resummed result can then be cast in various equivalent forms, and shown to sum to all orders in $\alpha_{s}(Q^{2})$ contributions corresponding to a given fixed logarithmic accuracy by including a suitable finite number of terms in the expansion of the coefficients that enter its expression.
@@ 7.1 The soft scale and the factorization scale
$C\left(N,\frac{Q^{2}}{\mu^{2}_{F}},\alpha_{s}(Q^{2})\right)=C^{(c)}\left(\frac{Q^{2}}{\mu_{F}^{2}},\alpha_{s}(Q^{2})\right)C^{(l)}\left(\frac{Q^{2}/(N)}{\mu^{2}_{F}},\alpha_{s}(Q^{2})\right)+O\left(\frac{1}{N}\right)$ (95)
@@
$Q^{2}\frac{d}{dQ^{2}}C^{(c)}\left(\frac{Q^{2}}{\mu^{2}_{F}},\alpha_{s}(Q^{2})\right)$ $=\bar{g}_{0}(\alpha_{s}(Q^{2}))C^{(c)}\left(\frac{Q^{2}}{\mu^{2}_{F}},\alpha_{s}(Q^{2})\right)$ (96)
$Q^{2}\frac{d}{dQ^{2}}C^{(l)}\left(\frac{Q^{2}/N}{\mu^{2}_{F}},\alpha_{s}(Q^{2})\right)$ $=\left[\int_{Q^{2}}^{Q^{2}/N^{2}}\frac{d\lambda^{2}}{\lambda^{2}}\bar{g}(\alpha_{s}(\lambda^{2}))\right]C^{(l)}\left(\frac{Q^{2}/N}{\mu^{2}_{F}},\alpha_{s}(Q^{2})\right).$ (97)
@@
$\ln C^{(l)}\left(\frac{Q^{2}/N^{2}}{\mu^{2}},\alpha_{s}(\mu^{2})\right)-\ln C^{(l)}\left(\frac{Q_{0}^{2}/N^{2}}{\mu^{2}},\alpha_{s}(\mu^{2})\right)=\int_{1}^{N^{2}}\frac{dn}{n}\,\int_{Q_{0}^{2}/N^{2}}^{Q^{2}/N^{2}}\frac{dk^{2}}{k^{2}}\,\bar{g}\left(\alpha_{s}(k^{2})\right),$ (98)
@@
$g(\alpha_{s})=-2A(\alpha_{s})+\frac{\partial B(\alpha_{s}(k^{2}))}{\partial\ln k^{2}}$ (99)
@@
$C^{(l)}\left(\frac{Q^{2}/N^{2}}{\mu^{2}_{F}},\alpha_{s}(\mu_{F}^{2})\right)=\exp\left\{\int_{1}^{N^{2}}\frac{dn}{n}\left[\left(-\int_{\mu^{2}(\mu_{F})}^{Q^{2}/N^{2}}\frac{dk^{2}}{k^{2}}2A(\alpha_{s}(k^{2}))\right)+B(\alpha_{s}(Q^{2}/n))\right]\right\},$ (100)
where $\mu^{2}(\mu_{F})$ is some reference scale that depends on $\mu_{F}$, and again we have introduced a factor of two as in Eq. (63) in order to keep into account the fact that there are two incoming partons.
We can determine both the value of the scale $\mu$ and of the function $A(\alpha_{s}(k^{2}/n))$ in Eq. (100) by demanding that the $\mu_{F}$ dependence of $C^{(l)}$ be given by the renormalization group equation (64), as we now show. First, we note that Eq. (40), together with the Mellin transformation formula Eq. (75), implies that $\gamma_{qq}$ is proportional to $\ln N$, up to constants and terms that vanish as $N\to\infty$. This turns out to be the case to all perturbative orders (in the $\overline{\rm MS}$ factorization scheme):
@@
The coefficient of the logarithmically enhanced contribution to the anomalous dimension is known as cusp anomalous dimension. This logarithmically enhanced contribution is present in the $\gamma_{qq}$ and $\gamma_{gg}$ anomalous dimension, that respectively characterize collinear emission of a gluon from a quark or a gluon line. Collinear emission of a quark from a gluon or a quark is not logarithmically enhanced, and in fact the corresponding anomalous dimensions vanish as $N\to\infty$. For the Drell-Yan process, on which we are focusing, the quark cusp anomalous dimension is relevant.
@@
$\mu_{F}^{2}\frac{d}{d\mu_{F}^{2}}\ln C^{(l)}\left(\frac{Q^{2}/N^{2}}{\mu_{F}^{2}},\alpha_{s}(\mu^{2})\right)=-2\ln NA(\alpha_{s}(\mu_{F}^{2})).$ (102)
Comparing to Eq. (64) with the anomalous dimension given by Eq. (101) (since $C^{(l)}$ only includes logarithmically enhanced contributions) we find that the resummation function $A(\alpha_{s}(\mu^{2}))$ coincides with the cusp anomalous dimension:
@@ 7.2 The resummed result and its accuracy
$C\left(N,\frac{Q^{2}}{\mu_{F}^{2}},\alpha_{s}(Q^{2})\right)=C^{(c)}\left(\alpha_{s}(Q^{2})\right)\exp\int_{1}^{N^{2}}\frac{dn}{n}\left[\left(-\int_{\mu_{F}^{2}}^{Q^{2}/n}\frac{dk^{2}}{k^{2}}2A(\alpha_{s}(k^{2}))\right)+B(\alpha_{s}(Q^{2}/n))\right],$ (104)
with $A$ given by the cusp anomalous dimension Eq. (103), and the functions $C^{(c)}$ and $B$ given as power series in $\alpha_{s}$:
@@
the coefficient of the double log Eq. (33), that now is exponentiated. Higher-order contributions to the cusp anomalous dimension correspond to double logs associated to higher-order corrections to the splitting function and anomalous dimension in this soft-collinear limit. Finally, the function $B(\alpha_{s})$ resums to all orders contributions that are associated to soft radiation, but not in the ordered collinear region of Sect. 4.2.
@@
log accuracy. accuracy of $C$: $\alpha_{s}^{n}L^{k}$ accuracy of $\ln C$: $\alpha_{s}^{n}L^{k}$ $g_{0}$: $\alpha_{s}^{k}$ $g_{j}$ order AA: $\alpha_{s}^{i}$ $B$: $\alpha_{s}^{i}$
@@
Subsequent contributions $g_{i}$ in the exponent correspond to the leading, net-to-leading,…, logarithmic approximation, where at the leading order the power of $\alpha_{s}$ is always by unit lower than the power of $\ln N$, at next-to-leading it is the same, at next-to-next-to-leading one less and so on. In the coefficient function at the leading logarithmic level each order in $\alpha_{s}$ is accompanied by an extra power of $\ln^{2}N$, and it is fully predicted by knowledge of $g_{1}$, which in turn is entirely predicted by knowledge of the $A_{1}$ coefficient, i.e. the leading-order cusp anomalous dimension. The function $g_{2}$ is predicted by knowledge of $A_{2}$ and $B_{1}$. Knowledge of $g_{2}$ then predicts at each order in $\alpha_{s}$ the coefficient of the next-lower order power of $\ln N$. However, if this is supplemented also by the first order coefficient $g^{1}_{0}$ in the expansion of the prefactor function $g_{0}$, at each order in $\alpha_{s}$ the coefficients of two next lower powers of $\ln N$ are actually predicted. This is usually referred to as next-to-leading logarithmic (NLL) approximation in the QCD literature. Confusingly in the SCET literature it is usually called NLL’, while NLL is the approximation without the $g^{1}_{0}$ coefficient, in which the coefficient of one less power of $\ln N$ is predicted. The pattern continues at next orders, and it is summarized in Table 1.
@@
The coefficients $A_{k}$, $B_{k}$ and $C^{(c)}_{k}$ can be determined by matching the resummed result to a fixed-order computation. Specifically, the leading cusp anomalous dimension $A_{1}$ is determined by the coefficient of the highest power of log of a next-to-leading (NLO) order computation. Comparison to the NLO also determines the coefficients $B_{1}$ and $C^{(c)}_{1}$ that enter the next-to-leading logarithmic (NLL) resummation. However, the NLO cusp anomalous dimension $A_{2}$ only appears in a NNLO fixed order computation. It is easy to check that this pattern persists to all orders: the coefficients $B_{k}$ and $C^{(c)}_{k}$ that enter the NkLL resummation are fully determined by a fixed NkLO fixed order result, but the NkLO cusp anomalous dimension only appears in a Nk+1LO fixed order computation. On the other hand, the cusp anomalous dimension is process-independent, so one the cusp anomalous dimension is known to NkLO, a NkLO fixed order computation of a process fully determines the coefficients needed for NkLL resummation.
@@
The resummed result Eq. (104) takes an especially simple form when choosing a factorization scale that coincides with the hard scale, $\mu_{F}^{2}=Q^{2}$. With this choice the exponential evolution factor is absent in the expression Eq. (67) of the coefficient function, because it is entirely reabsorbed in the parton luminosity, so the coefficient function becomes a function of $N$ and $\alpha_{s}(Q^{2})$ only. In this case its resummed expression becomes
$C(N,1,\alpha_{s}(Q^{2}))$ $=C^{(c)}\left(\alpha_{s}(Q^{2})\right)\exp\int_{1}^{N^{2}}\frac{dn}{n}\left[\left(-\int_{Q^{2}}^{Q^{2}/n}\frac{dk^{2}}{k^{2}}2A(\alpha_{s}(k^{2}))\right)+B(\alpha_{s}(Q^{2}/n))\right].$ (110)
@@
$C(N,1,\alpha_{s}(Q^{2}))$ $=C^{(c)}\left(\alpha_{s}(Q^{2})\right)\exp\int_{Q^{2}}^{Q^{2}/N}\frac{dk^{2}}{k^{2}}\left[2A(\alpha_{s}(k^{2}))\int_{N^{2}}^{Q^{2}/k^{2}}\frac{dn}{n}+B(\alpha_{s}(k^{2}))\right]$
$=C^{(c)}\left(\alpha_{s}(Q^{2})\right)\exp\int_{Q^{2}}^{Q^{2}/N^{2}}\frac{dk^{2}}{k^{2}}\left[2A(\alpha_{s}(k^{2}))\ln\frac{Q^{2}/N^{2}}{k^{2}}+B(\alpha_{s}(k^{2}))\right].$ (111)
This form of the resummed result shows that the $A$ term effectively resums logs by performing a further perturbative evolution with the large $N$ anomalous dimension Eq. (101) from the hard scale $Q^{2}$ to the soft scale $Q^{2}/N$.
@@
$\int_{0}^{1}dx\,\frac{x^{N-1}-1}{1-x}\ln^{k}(1-x)=-\sum_{k=0}^{\infty}\frac{\Gamma^{(k)}(1)}{k!}\frac{d^{k}}{dL^{k}}\int_{0}^{1-1/N}\frac{dx}{1-x}\ln^{k}(1-x)+O(1/N).$ (112)
Using this in Eq. (7.2) we get
@@
$\times$ $\exp\left[\int_{0}^{1}\!dx\frac{x^{N-1}-1}{1-x}\left(\int_{M^{2}}^{\left(1-z\right)^{2}Q^{2}}\frac{dk^{2}}{k^{2}}2A(\alpha_{s}(k^{2})\right)+\bar{B}(\alpha_{s}((1-x)^{2}Q^{2}))\right],$ (113)
where the function $\bar{B}$ is a series in $\alpha_{s}$ whose coefficients are determined order by order using Eq. (112) in terms of the coefficients $A_{k}$ and $B_{k}$: for instance $\bar{B}_{1}=B_{1}+2\gamma_{E}A_{1}$ and so on. This is the form of the resummation that was given in the original paper Eq. [Catani:1989ne]. The term proportional to $A$ in the exponent is just the Mellin transform of $P_{qq}(\alpha_{s}(k^{2}))$, integrated from scale $Q^{2}$ to scale $Q^{2}/N$, thereby exposing the physical meaning of the resummation. On the other hand, it should be noted that the integral over $k^{2}$ in Eq. (7.2) is ill-defined, because the strong coupling $\alpha_{s}(k^{2})$ diverges at the so-called Landau pole, i.e. when $k^{2}$ is small enough that the denominator in Eq. (70) vanishes. Hence, Eq. (7.2) is only meaningful insofar as it is equal to Eq. (7.2) up to terms that vanish in the $N\to\infty$ limit.

### v2

@@ 7 The resummed coefficient function
The resummed result given in Eq. (6.3) provides an expression for the physical anomalous dimension, i.e. essentially for the scale dependence of the coefficient function. Hence, a little more work is required in order to arrive at a resummed expression for the coefficient function itself. This is done by separating off the part of the $Q^{2}$ dependence of the coefficient function that goes through the ratio to the factorization scale $Q^{2}/\mu_{F}^{2}$. The final resummed result can then be cast in various equivalent forms, and shown to sum to all orders in $\alpha_{s}(Q^{2})$ contributions corresponding to a given fixed logarithmic accuracy by including a suitable finite number of terms in the expansion of the coefficients that enter its expression.
@@ 7.1 The soft scale and the factorization scale
$C\left(N,\frac{Q^{2}}{\mu^{2}_{F}},\alpha_{s}(Q^{2})\right)=C^{(c)}\left(\alpha_{s}(Q^{2})\right)C^{(l)}\left(\frac{Q^{2}/N^{2}}{\mu^{2}_{F}},\alpha_{s}(Q^{2})\right)+O\left(\frac{1}{N}\right)$ (95)
@@
$Q^{2}\frac{d}{dQ^{2}}C^{(c)}\left(\alpha_{s}(Q^{2})\right)$ $=\bar{g}_{0}(\alpha_{s}(Q^{2}))C^{(c)}\left(\alpha_{s}(Q^{2})\right)$ (96)
$Q^{2}\frac{\partial}{\partial Q^{2}}C^{(l)}\left(\frac{Q^{2}/N^{2}}{\mu^{2}_{F}},\alpha_{s}(Q^{2})\right)$ $=\left[\int_{Q^{2}}^{Q^{2}/N^{2}}\frac{d\lambda^{2}}{\lambda^{2}}\bar{g}(\alpha_{s}(\lambda^{2}))\right]C^{(l)}\left(\frac{Q^{2}/N^{2}}{\mu^{2}_{F}},\alpha_{s}(Q^{2})\right).$ (97)
@@
$\ln C^{(l)}\left(\frac{Q^{2}/N^{2}}{\mu^{2}},\alpha_{s}(Q^{2})\right)-\ln C^{(l)}\left(\frac{Q_{0}^{2}/N^{2}}{\mu^{2}},\alpha_{s}(Q_{0}^{2})\right)$ $=\int_{Q_{0}^{2}}^{Q^{2}}\frac{dk^{2}}{k^{2}}\int_{k^{2}}^{k^{2}/N^{2}}\frac{d\lambda^{2}}{\lambda^{2}}\bar{g}(\alpha_{s}(\lambda^{2}))$
$=-\int_{1}^{N^{2}}\frac{dn}{n}\,\int_{Q_{0}^{2}/n}^{Q^{2}/n}\frac{dk^{2}}{k^{2}}\,\bar{g}\left(\alpha_{s}(k^{2})\right),$ (98)
@@
$\bar{g}(\alpha_{s})=A(\alpha_{s})-\frac{\partial B(\alpha_{s}(k^{2}))}{\partial\ln k^{2}}$ (99)
@@
$C^{(l)}\left(\frac{Q^{2}/N^{2}}{\mu^{2}_{F}},\alpha_{s}(Q^{2})\right)=\exp\left\{\int_{1}^{N^{2}}\frac{dn}{n}\left[\left(-\int_{\mu^{2}(\mu_{F})}^{Q^{2}/N^{2}}\frac{dk^{2}}{k^{2}}A(\alpha_{s}(k^{2}))\right)+B(\alpha_{s}(Q^{2}/n))\right]\right\},$ (100)
where $\mu^{2}(\mu_{F})$ is some reference scale that depends on $\mu_{F}$.
We can determine both the value of the scale $\mu$ and of the function $A(\alpha_{s}(k^{2}))$ in Eq. (100) by demanding that the $\mu_{F}$ dependence of $C^{(l)}$ be given by the renormalization group equation (65), as we now show. First, we note that Eq. (40), together with the Mellin transformation formula (75), implies that $\gamma_{qq}$ is proportional to $\ln N$, up to constants and terms that vanish as $N\to\infty$. This turns out to be the case to all perturbative orders (in the $\overline{\rm MS}$ factorization scheme):
@@
The coefficient of the logarithmically enhanced contribution to the anomalous dimension is known as cusp anomalous dimension. This logarithmically enhanced contribution is present in the $\gamma_{qq}$ and $\gamma_{gg}$ anomalous dimensions, that respectively characterize collinear emission of a gluon from a quark or a gluon line. Collinear emission of a quark from a gluon or a quark is not logarithmically enhanced, and in fact the corresponding anomalous dimensions vanish as $N\to\infty$. For the Drell-Yan process, on which we are focusing, only the quark cusp anomalous dimension is thus relevant.
@@
$\mu_{F}^{2}\frac{d}{d\mu_{F}^{2}}\ln C^{(l)}\left(\frac{Q^{2}/N^{2}}{\mu_{F}^{2}},\alpha_{s}(Q^{2})\right)=\ln N^{2}A(\alpha_{s}(\mu_{F}^{2})).$ (102)
Note the factor of 2 from $\ln N^{2}=2\ln N$ exactly provides the factor of 2 needed to obtain twice $\gamma_{qq}$ as in Eq. (64). Comparing to Eq. (65) with the anomalous dimension given by Eq. (101) (since $C^{(l)}$ only includes logarithmically enhanced contributions) we find that the resummation function $A(\alpha_{s}(\mu^{2}))$ coincides with the cusp anomalous dimension:
@@ 7.2 The resummed result and its accuracy
$C\left(N,\frac{Q^{2}}{\mu_{F}^{2}},\alpha_{s}(Q^{2})\right)=C^{(c)}\left(\alpha_{s}(Q^{2})\right)\exp\int_{1}^{N^{2}}\frac{dn}{n}\left[\left(-\int_{\mu_{F}^{2}}^{Q^{2}/n}\frac{dk^{2}}{k^{2}}A(\alpha_{s}(k^{2}))\right)+B(\alpha_{s}(Q^{2}/n))\right],$ (104)
with $A$ given by the cusp anomalous dimension of Eq. (103), and the functions $C^{(c)}$ and $B$ given as power series in $\alpha_{s}$:
@@
the coefficient of the double log Eq. (33), that now is exponentiated. Higher-order contributions to the cusp anomalous dimension correspond to double logs associated to higher-order corrections to the splitting function and the anomalous dimension in this soft-collinear limit. Finally, the function $B(\alpha_{s})$ resums to all orders contributions that are associated to soft radiation, but not in the ordered collinear region of Sect. 4.2.
@@
log accuracy accuracy of $C$: accuracy of $\ln C$: $g_{0}$: $g_{j}$ AA: $B$:
$\alpha_{s}^{n}L^{k}$ $\alpha_{s}^{n}L^{k}$ $\alpha_{s}^{k}$ order $\alpha_{s}^{i}$ $\alpha_{s}^{i}$
@@
Subsequent contributions $g_{i}$ in the exponent correspond to the leading, next-to-leading,…, logarithmic approximation, where at the leading order the power of $\alpha_{s}$ is always by unit lower than the power of $\ln N$, at next-to-leading it is the same, at next-to-next-to-leading one more and so on. In the coefficient function at the leading logarithmic level each order in $\alpha_{s}$ is accompanied by an extra power of $\ln^{2}N$, and it is fully predicted by knowledge of $g_{1}$, which in turn is entirely predicted by knowledge of the $A_{1}$ coefficient, i.e. the leading-order cusp anomalous dimension. The function $g_{2}$ is predicted by knowledge of $A_{2}$ and $B_{1}$. Knowledge of $g_{2}$ then predicts at each order in $\alpha_{s}$ the coefficient of the next-lower order power of $\ln N$. However, if this is supplemented also by the first order coefficient $g^{1}_{0}$ in the expansion of the prefactor function $g_{0}$, at each order in $\alpha_{s}$ the coefficients of two next lower powers of $\ln N$ are actually predicted. This is usually referred to as next-to-leading logarithmic (NLL) approximation in the QCD literature. Confusingly in the SCET literature it is usually called NLL’, while NLL is the approximation without the $g^{1}_{0}$ coefficient, in which the coefficient of one less power of $\ln N$ is predicted. The pattern continues at next orders, and it is summarized in Table 1.
@@
The coefficients $A_{k}$, $B_{k}$ and $C^{(c)}_{k}$ can be determined by matching the resummed result to a fixed-order computation. Specifically, the leading cusp anomalous dimension $A_{1}$ is determined by the coefficient of the highest power of log of a next-to-leading (NLO) order computation. Comparison to the NLO also determines the coefficients $B_{1}$ and $C^{(c)}_{1}$ that enter the next-to-leading logarithmic (NLL) resummation. However, the NLO cusp anomalous dimension $A_{2}$ only appears in a NNLO fixed order computation. It is easy to check that this pattern persists to all orders: the coefficients $B_{k}$ and $C^{(c)}_{k}$ that enter the NkLL resummation are fully determined by a fixed NkLO fixed order result, but the NkLO cusp anomalous dimension only appears in a Nk+1LO fixed order computation. On the other hand, the cusp anomalous dimension is process-independent, so once the cusp anomalous dimension is known to NkLO, a NkLO fixed order computation of a process fully determines the coefficients needed for NkLL resummation.
@@
The resummed result of Eq. (104) takes an especially simple form when choosing a factorization scale that coincides with the hard scale, $\mu_{F}^{2}=Q^{2}$. With this choice the exponential evolution factor is absent in the expression (68) of the coefficient function, because it is entirely reabsorbed in the parton luminosity, so the coefficient function becomes a function of $N$ and $\alpha_{s}(Q^{2})$ only. In this case its resummed expression becomes
$C(N,1,\alpha_{s}(Q^{2}))$ $=C^{(c)}\left(\alpha_{s}(Q^{2})\right)\exp\int_{1}^{N^{2}}\frac{dn}{n}\left[\left(-\int_{Q^{2}}^{Q^{2}/n}\frac{dk^{2}}{k^{2}}A(\alpha_{s}(k^{2}))\right)+B(\alpha_{s}(Q^{2}/n))\right].$ (110)
@@
$C(N,1,\alpha_{s}(Q^{2}))$ $=C^{(c)}\left(\alpha_{s}(Q^{2})\right)\exp\int_{Q^{2}}^{Q^{2}/N^{2}}\frac{dk^{2}}{k^{2}}\left[A(\alpha_{s}(k^{2}))\int_{N^{2}}^{Q^{2}/k^{2}}\frac{dn}{n}+B(\alpha_{s}(k^{2}))\right]$ (111)
$=C^{(c)}\left(\alpha_{s}(Q^{2})\right)\exp\int_{Q^{2}}^{Q^{2}/N^{2}}\frac{dk^{2}}{k^{2}}\left[A(\alpha_{s}(k^{2}))\ln\frac{Q^{2}/N^{2}}{k^{2}}+B(\alpha_{s}(k^{2}))\right].$ (112)
This form of the resummed result shows that the $A$ term effectively resums logs by performing a further perturbative evolution with the large $N$ anomalous dimension given by Eq. (101) from the hard scale $Q^{2}$ to the soft scale $Q^{2}/N^{2}$.
@@
$\int_{0}^{1}dx\,\frac{x^{N-1}-1}{1-x}\ln^{k}(1-x)=-\sum_{n=0}^{\infty}\frac{\Gamma^{(n)}(1)}{n!}\frac{d^{n}}{dL^{n}}\int_{0}^{1-1/N}\frac{dx}{1-x}\ln^{k}(1-x)+O(1/N)$ (113)
where
$L=\ln\frac{1}{N}.$ (114)
Using this in Eq. (111) we get
@@
$\times$ $\exp\left[\int_{0}^{1}\!dx\,\frac{x^{N-1}-1}{1-x}\left(\int_{Q^{2}}^{\left(1-x\right)^{2}Q^{2}}\frac{dk^{2}}{k^{2}}2A(\alpha_{s}(k^{2})+\bar{B}(\alpha_{s}((1-x)^{2}Q^{2}))\right)\right],$ (115)
where the function $\bar{B}$ is a series in $\alpha_{s}$ whose coefficients are determined order by order using Eq. (113) in terms of the coefficients $A_{k}$ and $B_{k}$: for instance $\bar{B}_{1}=B_{1}+2\gamma_{E}A_{1}$ and so on. This is the form of the resummation that was given in the original paper [4]. The term proportional to $A$ in the exponent is just the Mellin transform of $P_{qq}(\alpha_{s}(k^{2}))$, integrated from scale $Q^{2}$ to scale $Q^{2}/N$, thereby exposing the physical meaning of the resummation. On the other hand, it should be noted that the integral over $k^{2}$ in Eq. (7.2) is ill-defined, because the strong coupling $\alpha_{s}(k^{2})$ diverges at the so-called Landau pole, i.e. when $k^{2}$ is small enough that the denominator in Eq. (53) vanishes. Hence, Eq. (7.2) is only meaningful insofar as it is equal to Eq. (111) up to terms that vanish in the $N\to\infty$ limit.